what is the theoretical yield of iron oxide when 55.65g of iron (ii) nitrate reacts with 12.75g of aluminum oxide?
1
Expert's answer
2019-12-03T08:27:13-0500

Reaction: 3Fe(NO3)2 + Al2O3 = 3FeO + 2Al(NO3)3

Amount of substance: n(Fe(NO3)2) = m/M = 55,65/180 = 0,309 mol

n(Al2O3) = 12,75/102 = 0,125 mol

So, the Al2O3 is in excess. Therefore, n(Fe(NO3)3)=n(FeO)

So, the yeild of FeO: m = n•M = 0,309•72 = 22,25 g.